If 3*tanx*tany=1 then show that 2cos (x+y)=cos(x-y)

Grade:11

1 Answers

Arun
6 years ago
3 sinx siny= cosx cosy
now,
2 sinx siny= cosx cosy –sinx siny
cos(x –y) –cos(x+y) = cos(x+y)
now cos(x –y)= 2 cos(x+y)
hence proved